A friction-free, weighted piston-cylinder device contains 2.0 kg of water in a mixed state at a pressure of 150kPa and a quality of 0.60. As it sits there, heat is lost until the mixture becomes entirely a saturated liquid. How much heat is lost in the process?

Thanks!
Answer #1

classmate Date Page # Using Table A-5 saturated water - at P = 150 kPa and x =0.60 hf = 467-13 hfo=2226 kJ/k 1 # hi hft x,